OpenStudy (anonymous):

Which equation would best help solve the following problem? The height of a triangle is 5 m less than its base. The area of the triangle is 42 m2. Find the length of the base.

OpenStudy (anonymous):

42 =x(x – 5) 42 =1/2 (x)(x-5) 42 =x^2– 5 42 =1/2(x^2-5)